Last year, a women's professional organization made two small-business loans totaling $26,000 to young women beginning their own businesses. The money was lent at 5% and 9% simple interest rates. If the annual income the organization received from these loans was $1,940, what was each loan amount?
Answer by Boreal(15235)

x=amount at 5%
26,000-x = amount at 9%. You can use x and y, but this is easier.
0.05x + 0.09(26,000-x)=1940
0.05x+2340-0.09x=1940
-0.04x=-400
x=10000
26000-x=13500
Check
0.05*10000=500
0.09*16000=1440
Total is 1940, units in dollars